OpenStudy (anonymous):

Algebra help please? Compare and contrast these four functions using complete sentences. Focus on teh slope and y-intercept first, and then on any additional properties of each function. I alread have some of it, but I need some help.

OpenStudy (anonymous):

Here are the four functions:

OpenStudy (anonymous):

give me answer chice

OpenStudy (anonymous):

I already have that f(x) = 2x + 5 and it gives me j(x) = 2x - 5. I think g(x) = 4x - 8 and h(x) = 2x - 1 but I'm not sure of those two are correct.

OpenStudy (anonymous):

It's an essay question, not multiple choice

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@Nnesha Do you think you can help?

Nnesha (nnesha):

for h(x) line is increasing or decreasing ?

OpenStudy (anonymous):

decreasing

OpenStudy (anonymous):

It's starting at the five and going over 1 and down 2. I don't know how to write that though.

Nnesha (nnesha):

so what does it means ?

Nnesha (nnesha):

slope would be what ?

OpenStudy (anonymous):

1/2?

Nnesha (nnesha):

nope if line is decreasing then slope would be negative

OpenStudy (anonymous):

Oh yeah, so it'd be -1/2?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

nope 2 is right slope is rise over run = 2/1 =2 it's decreasing so would be negative

OpenStudy (anonymous):

Oh okay. So the function would be h(x) = 5x - 2?

Nnesha (nnesha):

hmm y-intercept is a point where line intersect y-axis y=mx+b where m is slope b is y-intercept

OpenStudy (anonymous):

y = mx + b sooo h(x) = -2x + 5

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Okay, thank you. Now what about the g(x) function. Did I get that equation right?

Nnesha (nnesha):

nope slope and y-intercept both are wrong try again show ur work so i can find out the mistakes :=)

OpenStudy (anonymous):

Well I said it was g(x) = 4x - 8 because under the x column it is increasing by 4 and under the g(x) colume it is increasing by -8.

Nnesha (nnesha):

x isn't increasing by 4 would you add 4 with -2 to get 2 ?

OpenStudy (anonymous):

Yes... -2 + 4 = 2 then 2 + 4 is 6

OpenStudy (anonymous):

Even my calculator says that -2 + 4 is 2

Nnesha (nnesha):

hmm that's not a good way to find slope you should use slope formula

Nnesha (nnesha):

y-intercept isn't 8

OpenStudy (anonymous):

g(x) is increasing by -8 though. -1 - 8 = -9 and -9 - 8 = -17

Nnesha (nnesha):

that's not right slope isn't 4 use the slope formula \[\huge\rm \frac{ y_2 -y_1 }{ x_2 -x_2 }\]

Nnesha (nnesha):

wait a sec i see

OpenStudy (anonymous):

I don't have coordinates so I can't use that formula.

Nnesha (nnesha):

that's the difference of y's and x's y's =-8 x's = 4

Nnesha (nnesha):

so replace y_2 -y_1 = -8 and x_2 - x_1 =4 that would give you slope

OpenStudy (anonymous):

I don't understand what you're telling me to do.

Nnesha (nnesha):

 coordinates